OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Please show all of your work. Then graph the equation and submit your graph through the dropbox. -x – 3y =

OpenStudy (anonymous):

-x-3y=4

OpenStudy (mertsj):

Find the x intercept by replacing y with 0. Find the y intercept by replacing x with 0

OpenStudy (anonymous):

x-intercept is when y = 0, so \[-x -3(0) = 4 \rightarrow x = -4\] y intercept is when x = 0 \[-0 -3(y) = 4 \rightarrow y = -4/3\]

OpenStudy (mertsj):

So if y is 0, -x = 4 so x = -4 and the x intercept is (-4,0). If x is 0 -3y = 4 and y = -4/3 so the y intercept is (0,-4/3)

OpenStudy (anonymous):

y intercept = -3y/-3 = 4/-3 = - 4/3 = (0,- 4/3 x intercept = -x/-x= 4/-x = -4 = (-4,0)